OpenStudy (anonymous):

whats the sqarue root of 160 ?

OpenStudy (anonymous):

12.6 i think

OpenStudy (anonymous):

4sqrt10

OpenStudy (anonymous):

wait i did that wrong lol

OpenStudy (anonymous):

yes its 12.6

OpenStudy (taylev105):

12.64911 exactly

OpenStudy (anonymous):

no thats right

OpenStudy (arianna842):

12.6

OpenStudy (anonymous):

it is right

OpenStudy (anonymous):

\[\sqrt{160}=\sqrt{16*10}=\sqrt{16}*\sqrt{10}\] \[=4\sqrt{10}\]

OpenStudy (anonymous):

4sqrt10=12.6

OpenStudy (anonymous):

Taylev105 is right

OpenStudy (anonymous):

Thanks Guys!

OpenStudy (anonymous):

hey can someone help me with my question : http://openstudy.com/study#/updates/546a544de4b0db932291ac32